# ALiiCE: Evaluating Positional Fine-grained Citation Generation

### v1

@@ 3.4.2 Positional Fine-grained Citation Precision
Table 2: Results on ASQA when only outputs with positional fine-grained citations are evaluated. We omit the string "-psg" in the model settings for clarity. The best performances are highlighted in bold.
3.4.3 Coefficient of Variation of Citation Positions
Positional fine-grained citation generation allows citation marks to appear anywhere within a sentence, as long as their placement is logical and coherent. This means that the capability of generator $\mathcal{M}$ for positional fine-grained citation can, to some extent, be reflected from the degree of dispersion of citation marks positions. To quantify this capability, we propose CPCV (Coefficient of Variation of Citation Positions).
For response $\mathcal{R}$, we first calculate the citation position for every sentence. For sentence $s_{i}$, the position of citation marks can be expressed as:
$p_{i,{j}^{\prime}}=\frac{j\cdot\mathbb{1}_{s_{i,j}=\mathcal{C}_{i,j}}}{\left|s_{i}\right|}$ (3)
where ${j}^{\prime}$ is a new subscript because we omit words from $s_{i}$ in $p_{i}$. And $p_{i,{j}^{\prime}}$ ranges from 0 to 1. Then we can calculate the standard deviation of $s_{i}$ as:
[…]
where $\mu_{i}=\frac{1}{\left|p_{i}\right|}\sum^{\left|p_{i}\right|}_{{j}^{\prime}=1}p_{i,{j}^{\prime}}$, which represents the average value of $p_{i}$. Finally, we give the coefficient of variation of citation positions of response $\mathcal{R}$:
$CV_{CP}\left(\mathcal{R}\right)=\frac{1}{n}\sum^{n}_{i=1}\frac{\sigma\left(s_{i}\right)}{\mu_{i}}$ (5)
Obviously, a higher CPCV indicates a greater dispersion of citation positions. In practice, there may be sentences in $\mathcal{R}$ that do not have citation markers, thus we do not include these in practice.
Model ($k$-psg-form) ALiiCE ALCE Fluency Correct. CPCV Length
@@ 3.4.3 Coefficient of Variation of Citation Positions
GPT-3.5 (5-psg) 61.0 (0.5) 58.6 (2.2) 61.9 (0.9) 57.7 (0.6) 21.8 (0.6) 20.8 (0.3) 0.10 (-) 131.7 (46.2)
GPT-3.5 (5-psg-sum) 53.9 (2.6) 52.0 (1.1) 54.7 (2.3) 51.4 (2.5) 21.3 (5.1) 20.8 (1.1) 0.15 (-) 111.3 (46.5)
GPT-3.5 (5-psg-snippet) 53.4 (1.3) 50.9 (1.1) 53.6 (1.1) 50.8 (1.6) 34.9 (7.3) 20.8 (0.4) 0.13 (-) 106.7 (47.9)
GPT-3.5 (10-psg) 58.1 (2.4) 56.8 (2.0) 59.5 (1.7) 55.5 (1.0) 18.5 (4.7) 19.7 (0.7) 0.12 (-) 155.9 (57.4)
GPT-4 (5-psg) 55.1 (0.5) 54.0 (3.0) 57.9 (2.2) 53.2 (2.1) 20.4 (7.2) 21.3 (0.9) 0.15 (-) 102.2 (59.7)
LLaMA-3-8B (5-psg) 45.9 (0.3) 47.1 (0.7) 47.9 (2.1) 46.3 (1.0) 36.2 (1.0) 20.5 (0.9) 0.53 (-) 203.9 (71.4)
LLaMA-3-8B (10-psg) 42.8 (0.8) 44.2 (0.9) 43.6 (1.7) 42.4 (0.5) 32.5 (6.2) 19.5 (0.7) 0.61 (-) 224.2 (77.7)
Table 3: Results on ELI5. The correctness here refers to the exact match recall. Other descriptions follow Table 1.
@@
4 Experimental Setup
In this section, we describe the datasets and implementation details of our experiments.
4.1 Datasets
We utilize two popular datasets for the task of long-form QA, which are as follows:
ASQA
ELI5
The queries of these two datasets are well suited for retrieval-augmented generation, thus more conducive for evaluating fine-grained citation generation. Following (Gao et al., 2023b), we use the Generalizable T5-based dense Retriever  (GTR; Ni et al., 2022) to retrieve relevant passages for queries from Wikipedia corpus snapshot dated 2018-12-20.
4.2 Implementation
In our implementation, we utilize the SpaCy [^2] https://spacy.io/ to construct dependency trees for sentences, which is a useful and efficient python toolkit for many NLP tasks. We use TRUE [^3] https://huggingface.co/google/t5_xxl_true_nli_mixture, a fine-tuned T5-11B  (Raffel et al., 2020) model as the NLI model for the judgement of entailment in citation quality.
4.3 Models
We evaluate both closed-source and open-source LLMs on the task of positional fine-grained citation generation. For closed-source LLMs, we evaluate gpt-4-turbo-2024-04-09 and gpt-3.5-turbo-0125  (OpenAI, 2022; OpenAI et al., 2024). For open-source LLMs, we evaluate LLaMA-3-8B  (AI@Meta, 2024) and set top_p=0.95 for Nucleus Sampling  (Holtzman et al., 2020). We set the sampling temperature to 0.5 for all models. In addition, we also incorporate variables such as the number of retrieved documents and the document form used in generation (truncated original text, summary, or snippet) into the model setting. All prompts we employed are provided in Appendix C.
4.4 Evaluation Metrics
For comparison, we evaluate ALiiCE, representing the positional fine-grained citation evaluation method, and ALCE (Gao et al., 2023b), representing the sentence-level citation evaluation method.
In addition to the citation-specific metrics introduced above, we also utilize some common metrics in long-form QA. These metrics can also reflect potential influences caused by positional fine-grained citations, as detailed below:
Correctness
We check whether $\mathcal{R}$ answers the query $q$ accurately. For ASQA, we following (Stelmakh et al., 2022) to calculate exact match recall by checking whether ground truths are exact substrings of $\mathcal{R}$; For ELI5, we following (Fan et al., 2019) to use the F1 score of ROUGE-L.
Fluency
We quantify the fluency of text by MAUVE (Pillutla et al., 2021) to evaluate whether $\mathcal{R}$ generated by $\mathcal{M}$ is coherent, which is also essential for the task of long-form QA.
Length
We calculate the average length of $\mathcal{R}$ since it is sensitive to the long-form answers.
5 Main Results
The main results of our experiments on the ASQA and ELI5 datasets are presented in Table 1 and Table 3, respectively. In this section, we firstly present our key observations on the results, and then provide our detailed analyses on cases to show the shortcomings of sentence-level evaluation and the effectiveness of our approach.
5.1 Overall Performances
Difference in metric judgement.
There are differences in the evaluation on positional fine-grained citation text generation between ALiiCE and ALCE. Compared to ALCE, ALiiCE calculates lower citation recall, but higher citation precision. And this difference becomes more dramatic when only positional fine-grained citation outputs are evaluated, as illustrated in Table 2 and Table 4. We can observe this more intuitively in Figure 3. This change also leads to different best performing models under the two evaluation methods, according to the citation quality. GPT-3.5 with 5-passages setting performs best under the evaluation of ALCE, while ALiiCE tends to favor GPT-3.5 with 10-passages setting.
Difficulty in fine-grained generation.
We find that existing LLMs face challenges in generating positional fine-grained citations. LLMs still generate fewer fine-grained samples when prompts contain related documents. Although we employ In-Context Learning (ICL; Dong et al., 2023) to guide LLMs in placing citation marks by demonstration, this only produces slight improvement. The lower CPCV is indicative of this, which explains why ALCE and ALiiCE do not show very significant differences when evaluated on all citation texts. We suggest that positional fine-grained citation generation might be more suitable for specialized domains rather than open-domain QA.
Advances in open-source LLMs.
LLaMA-3 narrows the gap between open-source LLMs and closed-source LLMs in the citation text generation task. In previous studies, the citation quality of open-source LLMs is significantly worse than that of closed-source LLMs (Gao et al., 2023b; Huang et al., 2024). However, our experimental results show that the citation recall and precision of GPT-4 with 5-passages are only improved by 20.0% and 14.6%, respectively, compared to LLaMA-3-8B with 5-passages on ELI5. Additionally, LLaMA-3-8B has a higher CPCV and exhibits greater fluency, than both GPT-3.5 and GPT-4.
Figure 3: Comparison of citation recall and precision between ALCE and ALiiCE across three models on the ASQA dataset. All models use the 5-psg setting. The arrows indicate a uniform law: ALiiCE achieves a lower citation recall and higher citation precision.
Figure 4: Evaluation of citation quality by ALCE and ALiiCE on two examples from ASQA. The answers are generated by GPT-3.5 (5-psg). Case 1 demonstrates the difference between two methods on citation recall, and Case 2 illustrates the difference on citation precision.
5.2 Case Study
Long-context issue.
Citation precision issue.
There is a shortcoming in the calculation of citation precision in sentence-level evaluation, when assessing positional fine-grained citations. If there is an overlap between different evidences, it is potential for the NLI model to misjudge multiple atomic claims simultaneously. Taking the Case 2 in Figure 4 as an example, citation "[3]" contains evidences supporting both atomic claim 1 and 2. According to ALCE’s citation precision, citation "[3]" alone can support the entire sentence-level claim, whereas citation "[4]" cannot, as it only supports atomic claim 2. Consequently, citation "[4]" is considered redundant, despite it is actually a reasonable citation. In ALiiCE, we evaluate based on atomic claims, ensuring that the assessment is not influenced by evidences from other claims.
Based on the observed phenomenon of lower recall and higher precision, along with the previous analyses, we conclude that: compared to other metrics, ALiiCE has a higher decision threshold in the citation generation task. This indicates that ALiiCE is more conservative, only considering a citation correct when it has a high level of confidence. This is more beneficial for the citation generation task because the higher decision threshold encourages more accurate and relevant citations, reducing the likelihood of misleading information, which is particularly crucial in professional and high-risk fields (such as law and medicine) where incorrect citations can lead to serious consequences. Therefore, ALiiCE can help further enhance LLMs’ credibility while enabling effective evaluation.
Dataset Num of Claims Num of same NLI
ASQA 1935 1930
ELI5 3923 3891
5.3 Error Analyses
Grammatical error.
Parsing error.
Therefore, we conduct further experiment to test the potential impact of parsing errors on NLI. We firstly collect all the atomic claims from two datasets. Next, we utilize GPT-3.5 to refine each claim based on its original sentence (the prompt is provided in Appendix C). And then we employ the NLI model to assess the entailment before and after the claim refinement. As indicated in Table 5, the result show that the proportion of claims with inconsistent NLI results is less than 1% across both datasets. Therefore, the parsing error is unlikely to have a significant impact on the evaluation.
6 Related Work
Attribution.
Attribution refers to the ability of LMs to generate and provide evidence  (Li et al., 2023). The source of attribution can be pre-training data  (Han and Tsvetkov, 2022; Weller et al., 2024), or out-of-model knowledge  (Shuster et al., 2021; Li et al., 2024). When the source is documents, citation is a common form of attribution (Kamalloo et al., 2023).  Ye et al. (2024); Huang et al. (2024) study generating response and citations simultaneously, while Gao et al. (2023a); Huo et al. (2023) research on adding citations in the post-hoc stage.
Retrieval-Augmented Generation.
Citation Evaluation.
The current citation evaluation methods are mainly performed by human evaluation, which is costly and time-intensive (Chen et al., 2023). Subsequently, automatic evaluation methods are proposed, including classification-based metrics (Liu et al., 2023; Yue et al., 2023) and quantitative metrics (Gao et al., 2023b; Li et al., 2024). However, these methods are primarily sentence-level, leading to issues with atomicity of claims (Malaviya et al., 2024) and ambiguity (Liu et al., 2023). We propose ALiiCE, the first evaluation method for positional fine-grained citations.
7 Conclusion
We evaluate different LLMs and obtain several observations: 1) ALiiCE shows a higher decision threshold compared to sentence-level evaluation; 2) existing LLMs face challenges in generating positional fine-grained citations; 3) open-source LLMs narrow the gap in citation generation with closed-source LLMs. Our further analyses demonstrate insufficiency of existing sentence-level evaluation methods and effectiveness of our ALiiCE for assessing positional fine-grained citations. We hope that our work can inspire more research into positional fine-grained citation text generation.
Limitations
In our experiments, we only utilize the open-domain long-form QA datasets. However, positional fine-grained citation generation is applicable to a broader range of scenarios, such as academic writing and automatic summarization. Additionally, positional fine-grained citations are more likely to emerge when there are clear logical relationships between claims, such as multi-hop reasoning. Therefore, it is necessary to expand the data domain of the benchmark.
Ethics Statement
Appendix A Parsing Algorithm Details
  • If there is a subtree between $T_{i}$ and $T_{j}$ with a dependency relation of c̈cb̈etween its root node and the LCA node (we refer to this subtree $T_{c}$), then we discuss:
Relation Type Explanation
acomp adjectival complement
advcl adverbial clause modifier
amod adjectival modifier
cc coordination
conj conjunct
mark marker
nmod nominal modifier
nsubj nominal subject
nsubjpass passive nominal subject
pobj object of a preposition
prep prepositional modifier
punct punctuation
relcl adnominal relative clause modifier
Table 6: Several common types of dependency relation.
Appendix B Parsing Examples
To improve the intuitiveness of the parsing algorithm, we present three straightforward examples (Figures 5 to 11). Each figure shows a dependency tree, where each node represents a word node. For word nodes matched with citations (marked in red), the format of the node value is "word : index : citation marks", where "index" denotes the position of the word in the original sentence. For word nodes without citations (marked in green), the format of the node value is "word : index". The sentences to be parsed are all from the outputs of the GPT-3.5 (5-psg) on the ASQA dataset.

### v2

@@ 5.3 Error Analyses
Table 3: Results on ASQA when only outputs containing positional fine-grained citations are evaluated. We omit the string "-psg" in the model settings for clarity. The best performances are highlighted in bold.
6 Human Evaluation
We conduct human evaluation to examine the correlation between ALiiCE and human judgment. Since  Gao et al. (2023b); Liu et al. (2023) have thoroughly studied sentence-level citation evaluation, we only focus on LLMs’ responses that include positional fine-grained citations. In addition to the citation recall and precision, we also consider: 1) the proportion of positional fine-grained responses to total responses, 2) the answer utility, which assesses whether the LLM’s response is helpful in answering the question, and 3) the citation utility, evaluates whether the positional fine-grained citation is useful for the response. We recruit three annotators to evaluate the outputs of the models used in the previous experiments.
We observe that ALiiCE and human judgment show a strong correlation. The model rankings evaluated by ALiiCE align closely with those evaluated by human judgement. The average Cohen’s kappa coefficients between ALiiCE and annotators for ASQA are 0.71 for citation recall and 0.62 for citation precision, demonstrating high consistency. In addition, responses containing fine-grained citations constitute a small proportion of the total output. For instance, the fine-grained output of GPT-3.5 (5-psg-summ) on ELI5 accounts for only 8% of the total samples. This pattern is consistent with the results shown by CVCP. Details on the human evaluation are provided in the Appendix B.
7 Discussion
Figure 4: Comparison of citation recall and precision between ALCE and ALiiCE across three models using the 5-psg setting on ASQA. ALiiCE achieves lower citation recall and higher citation precision.
Based on the experimental results and observations, we discuss our insights on the task of positional fine-grained citation generation, as follows:
ALiiCE has a higher decision threshold.
Compared to ALCE, ALiiCE calculates lower citation recall, but higher citation precision. And this difference becomes more dramatic when only positional fine-grained citation outputs are evaluated, as illustrated in Table 3. We can observe this more intuitively in Figure 4. This means that ALiiCE has a higher decision threshold, indicating that ALiiCE is more conservative, only considering a citation correct when it has a high level of confidence. This is more beneficial for the citation generation task because the higher decision threshold encourages more accurate and relevant citations, reducing the likelihood of misleading information, which is particularly crucial in professional and high-risk fields (e.g., law and medicine) where incorrect citations can lead to serious consequences.
Open-source LLMs display great progress.
LLaMA-3 narrows the gap between open-source LLMs and closed-source LLMs in the citation text generation task. In previous studies, the citation quality of open-source LLMs is significantly worse than that of closed-source LLMs (Gao et al., 2023b; Huang et al., 2024a). However, our experimental results show that the citation recall and precision of GPT-4 with 5-passages are only improved by 20.0% and 14.6%, respectively, compared to LLaMA-3-8B with 5-passages on ELI5. Additionally, LLaMA-3-8B has a higher CVCP and exhibits greater fluency, than both GPT-3.5 and GPT-4.
Rethinking citation quality through the lens of citation utility.
Our human evaluation indicates that citation utility and citation quality do not show a strong correlation. And our annotators observe that in some responses, even when the citation utility score is zero, the citation quality remains high. Thus, existing citation quality metrics can only evaluate the correctness of citation marks for each claim, but they fail to assess the utility of these marks, as being correct is not equal to being useful. We believe that this is worth further exploration in future research on citation evaluation methods.
How to study positional fine-grained citation generation?
Through our observation of the fine-grained responses, we find that most atomic claims with sufficient citation utility, exhibit certain logical relationships, such as parallelism, causality, and transitions. Under these logical structures, positional fine-grained citations often have better utility and significantly enhance user-friendliness. Constructing reasoning paths for multi-step retrieval and generation can establish clearer logical relationships for long-form responses, thereby promoting fine-grained citations. Additionally, in the supervised learning method, creating labeled data presents a significant challenge.  Ye et al. (2024) design an algorithm for automatically annotating citation marks at sentence-level. However, this method becomes more challenging for positional fine-grained citation generation. Similarly, we recommend constructing supervised labels by multi-hop QA datasets and also combining sentence-level citation sequences to ensure generalization.
8 Related Work
Attribution
Attribution refers to the ability of LMs to generate and provide evidence  (Li et al., 2023). The source of attribution can be pre-training data  (Han and Tsvetkov, 2022; Weller et al., 2024), or out-of-model knowledge  (Shuster et al., 2021; Li et al., 2024). When the source is documents, citation is a common form of attribution (Kamalloo et al., 2023).  Ye et al. (2024); Huang et al. (2024a) study generating response and citations simultaneously, while Gao et al. (2023a); Huo et al. (2023) research on adding citations in the post-hoc stage.
Retrieval-Augmented Generation
Citation Evaluation
The current citation evaluation methods are mainly performed by human evaluation, which is costly and time-intensive (Chen et al., 2023). Thus, automatic evaluation methods are studied, including classification-based metrics (Liu et al., 2023; Yue et al., 2023) and quantitative metrics (Gao et al., 2023b; Li et al., 2024). In specific domains, Li et al. (2022); Li and Ouyang (2024) study the citation generation for academic writing. However, most research is primarily sentence-level, leading to issues with atomicity of claims (Malaviya et al., 2024) and ambiguity (Liu et al., 2023). We propose ALiiCE, the first evaluation method for positional fine-grained citations.
9 Conclusion
We evaluate several LLMs and observe that currently, LLMs lack strong capabilities for generating fine-grained citations. We demonstrate the need of designing dedicated method for positional fine-grained citation evaluation and the effectiveness of ALiiCE in addressing this need. We also discuss some useful conclusions: 1) the latest open-source LLMs narrow the gap between them and closed-source LLMs in citation generation; 2) current metrics for citation quality lack consideration of citation utility; 3) the logical relationships between atomic claims can be considered when designing methods for positional fine-grained citation generation. We hope that our work can inspire more research into this underexplored task.
Limitations
In our experiments, we only utilize the open-domain long-form QA datasets. However, positional fine-grained citation generation is applicable to a broader range of scenarios, such as academic writing and summarization. Therefore, it is necessary to expand the data domain of the benchmark.
Ethics Statement
Appendix A Experimental Setup Details
A.1 Datasets
We utilize two datasets of long-form QA, and a detailed description of them is as follows:
ASQA
ELI5
A.2 Models
For LLaMA-3-8B, we set top_p=0.95 for Nucleus Sampling  (Holtzman et al., 2020). And we set the sampling temperature to 0.5 for all models.
A.3 Evaluation Metrics
Regarding correctness, for ASQA, we follow Stelmakh et al. (2022) to calculate exact match recall by checking whether ground truths are exact substrings of $\mathcal{R}$; for ELI5, we following Fan et al. (2019) to use the F1 score of ROUGE-L. Regarding fluency, we quantify the fluency of text by MAUVE (Pillutla et al., 2021).
Datasets Model ($k$-psg-form) PF Sample Answer Utility PF Citation Utility Human ALiiCE
@@ A.3 Evaluation Metrics
ASQA GPT-3.5 (5-psg) 71 (7.5%) 3.4 (0.58) 0.62 (0.78) 77.0 75.2 75.4 74.2
GPT-3.5 (5-psg-summ) 161 (17.0%) 3.2 (0.64) 0.50 (0.73) 70.1 69.8 73.9 72.4
GPT-3.5 (5-psg-snip) 190 (20.0%) 3.5 (0.57) 0.47 (0.68) 63.9 62.4 60.5 62.6
GPT-3.5 (10-psg) 129 (13.6%) 3.2 (0.47) 0.49 (0.72) 77.8 78.8 75.8 77.9
GPT-4 (5-psg) 140 (14.8%) 3.7 (0.50) 0.63 (0.69) 70.5 75.5 69.3 75.7
LLaMA-3-8B (5-psg) 286 (30.2%) 3.0 (0.49) 0.45 (0.66) 58.1 66.7 56.9 64.3
LLaMA-3-8B (10-psg) 252 (26.7%) 2.7 (0.44) 0.41 (0.63) 58.9 67.2 57.7 66.1
ELI5 GPT-3.5 (5-psg) 85 (8.5%) 3.1 (0.57) 0.43 (0.67) 41.4 49.4 40.1 50.0
GPT-3.5 (5-psg-summ) 80 (8.0%) 2.8 (0.56) 0.46 (0.63) 37.1 35.8 35.9 35.1
GPT-3.5 (5-psg-snip) 104 (10.4%) 2.9 (0.42) 0.51 (0.60) 40.9 39.7 39.6 39.1
GPT-3.5 (10-psg) 132 (13.2%) 2.7 (0.48) 0.46 (0.63) 42.7 46.6 44.2 48.1
GPT-4 (5-psg) 119 (11.9%) 3.3 (0.55) 0.50 (0.62) 41.0 45.7 40.5 46.2
LLaMA-3-8B (5-psg) 230 (23.0%) 2.8 (0.51) 0.36 (0.70) 40.4 46.0 41.1 44.0
LLaMA-3-8B (10-psg) 207 (20.7%) 2.4 (0.66) 0.33 (0.61) 39.6 47.2 41.8 47.7
Table 4: Human evaluation results on ASQA and ELI5. The value in bracket of PF Sample is the percentage of responses containing positional fine-grained citation to the total responses. The value in bracket of Answer Utility and PF Citation Utility is the Fleiss’ Kappa coefficient of three annotators. Every value of human evaluation metrics in the table is the average of the results from three annotators.
Appendix B Human Evaluation Details
We conduct human evaluation to examine the correlation between ALiiCE and human judgment. We manually inspect only those samples containing positional fine-grained citations in the model output, as these are aligned with our task requirements. We focus on five metrics, as detailed below:
PF Sample
This represents positional fine-grained sample, which is the quantity of responses containing positional fine-grained citations.
Answer Utility
Whether LLM’s response is helpful in answering the question. We employ a 1-5 Likert scale, corresponding to Strongly Disagree, Disagree, Neutral, Agree, and Strongly Agree.
PF Citation Utility
Whether the positional fine-grained citation in the response is useful. We employ a binary annotation rule. PF Citation Utility is 1 when all the following conditions are met: 1) the positions of the fine-grained citations are reasonable (i.e., each citation corresponds to a clear atomic claim); 2) the fine-grained citations improve readability and user verifiability, and reduce ambiguity compared to sentence-level citations (for details, see Section 1); and 3) there is no excessive or redundant citation. Otherwise, PF Citation Utility is 0. For example, "Filming began in late May 2015[3], and the movie was released on March 25, 2016[3]." contains redundant citations and does not improve readability or user verifiability. Therefore, its PF Citation Utility is annotated as 0.
Citation Recall
This is the human-calculated citation recall. The annotator extract atomic claims manually and judge whether the cited passages can entail the claim. Its calculation is consistent with the description in Section 3.3.1.
Citation Precision
This is the human-calculated citation precision. The annotator judge whether each citation is redundant. Its calculation is consistent with the description in Section 3.3.2.
We recruit three annotators who are highly familiar with NLP research and well-acquainted with our work. The results of ASQA and ELI5 are shown in Table 4. Our analysis of the results is as follows:
ALiiCE and human judgement show a strong correlation.
We observe that the ranking of models evaluated by ALiiCE is consistent with the ranking based on human judgment. Furthermore, we calculate the Cohen’s kappa coefficient between ALiiCE and each annotator’s judgement, and the average result shows that the coefficient of citation recall is 0.71, and the coefficient of citation precision is 0.62, demonstrating high consistency. Additionally, under positional fine-grained citations, human judgment does not align with ALCE’s evaluation. This underscores the necessity of positional fine-grained evaluation methods, which cannot be substituted by sentence-level evaluation methods.
Current LLMs generate limited outputs containing positional fine-grained citation.
This indicates that LLMs still face difficulties in providing positional fine-grained citations, which echoes the observation in Section 5.1.
No direct relationship between citation utility and answer utility.
We suggests that in long-form QA, fine-grained citations often occur within supplementary explanations rather than in the core sentences of the answers. However, answer utility is mainly contributed by the core sentence. Hence, there is no strong correlation between them.
Citation utility should be given serious consideration.
This conclusion is consistent with that in Section 7. Our findings indicate that PF citation utility and citation quality do not demonstrate a strong correlation. Our annotators observed that in some samples, even when the utility score is zero, the citation quality remains high. For example, the PF Citation Utility of "Filming began in late May 2015[3], and the movie was released on March 25, 2016[3]." is 0, but the citation recall and citation precision are all 1. Therefore, existing citation quality metrics can only evaluate the correctness of citation marks for each claim, but they fail to assess the utility of these marks.
Appendix C Parsing Algorithm Details
  • If there is a subtree between $T_{i}$ and $T_{j}$ with a dependency relation of "cc" between its root node and the LCA node (we refer to this subtree $T_{c}$), then we discuss:
Relation Type Explanation
acomp adjectival complement
advcl adverbial clause modifier
amod adjectival modifier
cc coordination
conj conjunct
nmod nominal modifier
nsubj nominal subject
nsubjpass passive nominal subject
pobj object of a preposition
prep prepositional modifier
punct punctuation
Table 5: Several common types of dependency relation.
Appendix D Parsing Examples
To improve the intuitiveness of the parsing algorithm, we present three straightforward examples (Figures 5 to 13). Each figure shows a dependency tree, where each node represents a word node. For word nodes matched with citations (marked in red), the format of the node value is "word : index : citation marks", where "index" denotes the position of the word in the original sentence. For word nodes without citations (marked in green), the format of the node value is "word : index". The sentences to be parsed are all from the outputs of the GPT-3.5 (5-psg) on the ASQA dataset.
Appendix E Prompts
Appendix F CVCP Details
In Appendix B, we preliminarily verify the consistency of CVCP with the degree of positional fine-grained citations. In this section, we further analyze the meaning of the CVCP value and provide a reference. We use the output of GPT-3.5, GPT-4, and LLaMA-3-8B with 5-psg generated from ASQA. We randomly select 200 responses containing fine-grained citations (denoted as $E$) and 200 responses without fine-grained citations (denoted as $F$). The CVCP for $E$ is 0.85, while for $F$ it was 0. We then randomly select 100 samples from each of $E$ and $F$ to form $G$, and repeat the calculation five times, resulting that the average CVCP of $G$ is 0.67. The reference of CVCP here is not entirely sufficient, as it would be more reasonable to use gold answers written by human experts. Thus, it is necessary to design a dedicated datasets for long-form QA with positional fine-grained citations, which should be addressed in future work.
Appendix G Additional Results
We compare the evaluation results from ALCE and ALiiCE. Compared to ALCE, ALiiCE calculates lower citation recall, but higher citation precision. We present this difference at Table 3 and 6. This change also leads to different best models under the two evaluation methods, in terms of citation quality. GPT-3.5 with 5-passages setting performs best under the evaluation of ALCE, while ALiiCE tends to favor GPT-3.5 with 10-passages setting.